# DuSEGO: Dual Second-order Equivariant Graph Ordinary Differential Equation

### v1

@@ 5. DuSEGO Properties Analysis
where $\hat{D}=\max\limits_{i,j\in\mathcal{V}}\frac{1}{\sqrt{d_{i}d_{j}}}$, and $\Gamma_{x}:=6+4\max\limits_{x}|\sigma^{{}^{\prime}}(x)|\cdot\hat{D}\max\limits_{1\leq n\leq T}||\mathbf{w}^{n}||_{1}$ and $\Gamma_{h}:=6+4\max\limits_{h}|\sigma^{{}^{\prime}}(h)|\cdot\hat{D}\max\limits_{1\leq n\leq T}||\mathbf{w}^{n}||_{1}$. $d_{i}$ is the degree of node $i$, $\mathbf{\bar{X}}_{i}$ is the label of node $i$ and $\mathbf{\bar{H}}_{i}$ is the coordinate of node $i$.
Proof: From (rusch2022graph, ), we have:
@@
6. Experiments
6.1. Modelling a Dynamic System
In this section, we evaluate DuSEGO on one simulated N-body system named charged particles, where we aim to forecast the position of particles across different time steps (satorras2021n, ).
6.1.1. Dataset
Building on the work of (fuchs2020se, ), (satorras2021n, ) extended the Charged Particles N-body experiment introduced in (kipf2018neural, ), where the system consists of five particles, each defined by charge, position, and velocity in three-dimensional space. In this setup, particles in each graph are fully connected, with the initial velocity norm included as an additional node feature. Using the N-body system code from (satorras2021n, ), we generated a new dataset for the charged N-body system across 1,000, 1,500, and 2,000 timesteps. The initial positions of particles were sampled from a Gaussian distribution with $\mu=0$ and $\sigma=0.5$, and the initial velocity was represented as a random vector with a norm of 0.5. The dataset is split into training, validation, and testing sets, containing 10,000, 2,000, and 2,000 samples, respectively.
Table 1. Mean squared error (×$\left(\times 10^{-2}\right)$ for the future position estimation under different timesteps in the N-body system experiment, and forward time in seconds for a batch size of 100 samples running on NVIDIA A100 GPUs. Results averaged across 5 runs. We report both mean and standard deviation of MSE in the table. The best performance is bold.
Methods MSE Forward Time/ms
@@ 6.1.1. Dataset
6.1.2. Baselines.
Two types of baseline methods are deployed for comparison: (1) Methods that incorporate first-order information, including the Linear model, Graph Neural Networks (GNN), SE(3) Transformer (fuchs2020se, ), Tensor Field Networks (TFN) (thomas2018tensor, ), Radial Field (RF) (kohler2019equivariant, ), Equivariant Graph Neural Networks (EGNN) (satorras2021n, ), and Steerable E(3) Equivariant Graph Neural Networks (SEGNN) (brandstetter2021geometric, ); (2) Models that incorporate second-order information, including GraphCon (rusch2022graph, ), Graph Mechanics Network (GMN) (huang2021equivariant, ), and Second-order Equivariant Graph Neural Ordinary Differential Equation (SEGNO) (liu2023improving, ).
6.1.3. Experimental Setting.
We build on the experimental framework introduced in (liu2023improving, ), aiming to estimate the positions of all particles after 1,000, 1,500, and 2,000 timesteps. We re-evaluate all baseline methods under identical experimental settings on NVIDIA A100 GPUs, gathering results across five runs to ensure fair comparisons. The experimental setup is as follows: Adam optimizer with a learning rate of $5\times 10^{-4}$, batch size of 100, 10,000 epochs, hidden embedding dimension of 64, weight decay of $1\times 10^{-12}$, and four GNN layers. Specifically, for the GMN method, we set the numbers of isolated particles, sticks, and hinges to $(5,0,0)$. For the SE(3)-transformer and TFN, we set the representation degrees to 3 and 2, respectively. An early stopping criterion is applied, halting training if no improvement in validation loss is observed within 50 epochs.
6.1.4. Implementation details.
@@
(b) Dirichlet energy $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions.
Figure 4. Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features $\mathbf{H}^{n}$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node positions $\mathbf{X}^{n}$ propagated through a EGNN and DuSEGO-EGNN for two different values of $\alpha=0,1$ and $\gamma_{1}=\gamma_{2}=1$ in Eq. 8 and Eq. 9.
@@
The results for all baseline methods, DuSEGO-EGNN, and DuSEGO-SEGNN on the N-body system dataset are shown in Table 1. We observe that DuSEGO significantly enhances the performance of EGNN and SEGNN, with DuSEGO-SEGNN outperforming all other baseline methods and DuSEGO-EGNN under the same experimental conditions. Specifically, DuSEGO-EGNN and DuSEGO-SEGNN reduce the mean squared error (MSE) of SEGNN and EGNN by an average of 7.73% and 6.65%, respectively, across the three timesteps, indicating substantial improvement. Notably, the relative improvement for DuSEGO-EGNN increases from 5.86% at 1,000 timesteps to 13.13% at 1,500 timesteps, highlighting the strong generalization ability of DuSEGO. Additionally, Figure 2 presents a visualization of GMN, EGNN, SEGNN, GMN, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN in estimating particle positions for the N-body system experiment at 1,000 timesteps, with MSE values between predictions and ground truths reported. This visualization further shows that DuSEGO-SEGNN and DuSEGO-EGNN outperform EGNN and SEGNN, respectively, aligning with the performance results in Table 1.
@@
It is crucial to verify that the physical principles identified from a limited number of experiments are sufficiently general to explain phenomena on a broader scale. To this end, we assess the performance of EGNN, SEGNN, GMN, GraphCon, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN across various training set sizes, ranging from 500 to 3,000 samples in the N-body system dataset. In this evaluation, the timesteps of particle positions are set to 1,000. As shown in Fig. 3, DuSEGO-EGNN and DuSEGO-SEGNN, similar to the GMN method, demonstrate more stable performance than other baseline methods, emphasizing their data efficiency. This result suggests that data efficiency, a notable strength of DuSEGO, is essential for effectively learning to model physical systems.
@@
To further investigate the over-smoothing issue in EGNN, we implement EGNN and DuSEGO-EGNN with two different values of $\alpha=0,1$ and set $\gamma_{1}=\gamma_{2}=1$ in Eq. 8 and Eq. 9 on the N-body system dataset at 1,000 timesteps. As shown in Fig. 4, the over-smoothing problem in EGNN becomes apparent as the number of EGNN layers increases. The differences between node features, such as $\mathbf{h}_{i}^{n}$ and $\mathbf{h}_{j}^{n}$, as well as node positions, like $\mathbf{x}_{i}^{n}$ and $\mathbf{x}_{j}^{n}$, gradually decrease, indicating that node features and positions are converging toward a constant value. However, with the integration of DuSEGO into EGNN, the over-smoothing issue is effectively mitigated, as $\mathbf{E}\left(\mathbf{H}^{n}\right)$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ remain stable within a reasonable range, even as the number of EGNN layers increases.
@@
Additionally, we compare the performance of EGNN and DuSEGO-EGNN as the number of EGNN layers increases (up to 16). As shown in Table 2, the MSE for EGNN decreases from layer 4 to layer 8. However, at 12 layers, the gradient vanishing issue emerges, preventing EGNN from producing accurate results. This confirms that gradient vanishing becomes a limitation as the depth of EGNN increases. In contrast, the MSE for DuSEGO-EGNN continues to decrease with additional layers across different timesteps, and DuSEGO-EGNN maintains stable performance, supporting our hypothesis that integrating EGNN within DuSEGO helps mitigate the gradient vanishing problem.
@@
Table 2. Mean squared error (×$\left(\times 10^{-3}\right)$ for the future position estimation under different timesteps in the N-body system experiment between EGNN and DuSEGO-EGNN under different layers. Results averaged across 5 runs.
@@ 6.2. Graph Autoencoder
Subsequently, we compare DuSEGO with the state-of-art baseline methods for the Graph Autoencoder task, where we aim to estimate the adjacency matrix ${A}_{ij}$ of a graph $\mathcal{G}$ generated from the graph autoencoder and minimize the binary cross entropy between the estimated adjacency matrix $\hat{A}_{ij}$ and the ground truth $A_{ij}$.
@@
Two widely used datasets are deployed in this section: (1) Community  (kipf2016variational, ) dataset, generated using the Erdos-Renyi model (erdds1959random, ), consists of random graphs where nodes are connected with a predefined probability, capturing the diversity in the density and structure of various communities. This model, denoted as $G(n,p)$, generates networks by independently linking pairs of nodes with probability $p$; (2) Erdos & Renyi (janson2011random, ) dataset, based on the foundational random graph generation model (janson2011random, ; you2018graphrnn, ) and also represented as $G(n,p)$, includes a collection of graphs that exhibit the randomness and variability of network structures.
Following (satorras2021n, ), we generate the Community and Erdos & Renyi datasets (kohler2019equivariant, ; kipf2018neural, ) using the code provided in (kipf2018neural, ). Specially, we sample 50,000 graphs for training, 5,000 for validation, and 5,000 for testing, where we set $p=0.1$ and $6\leq n\leq 10$ for Erdos&Renyi, and $p=0.15$ and $6\leq n\leq 11$ for Community, respectively. Additionally, 20 communities are divided in the Community dataset, and each graph is defined as an adjacency matrix $A\in\{0,1\}^{M\times M}$ in these two datasets.
@@ 6.2.2. Baselines
Two types of baseline methods are deployed for comparison: (1) Methods that contains first-order information including GNN, GNN-Noise, RF (kohler2019equivariant, ), EGNN (satorras2021n, ); (2) Methods that contains second-order information including GraphCon (rusch2022graph, ), GMN (huang2021equivariant, ) and SEGNO (liu2023improving, ).
@@
The noise generated from $\mathcal{N}(\mathbf{0},\sigma\mathbf{I})\in\mathbb{R}^{M\times n}$ will be used as input for all baseline methods. The output will then be employed to embed the graph and fed into the graph decoder to generate an estimated adjacency matrix, $\hat{A}_{ij}$. We rerun all baseline methods under the same experimental settings on NVIDIA A100 GPUs and collect the results across five runs for a fair comparison. The experimental settings are as follows: Adam optimizer with learning rate ${1\times 10^{-3}}$, the number of epochs 1,000, batch size 256, hidden embedding dim 64, weight decay ${1\times 10^{-12}}$, layer number 4, and dimensions for the embedding space 8. Velocity is not considered in the Graph Autoencoder task, which is set to None in the EGNN, GMN, SEGNO and DuSEGO-EGNN.
@@
We build upon the experimental setting introduced in (satorras2021n, ). and employ EGNN as the backbone of DuSEGO, named DuSEGO-EGNN. F1 score of the edge classification and % Error, defined as the percentage of incorrectly predicted edges relative to the total number of potential edges, are reported here. EGNN is implemented based on the implementation details provided in (satorras2021n, ), which is used as the coupling function for message passing (rusch2022graph, ). Then, we use the ODE solvers provided in torchdiffeq library for solving Eq 8 and Eq. 9.
@@
Table 3 depicts the overall results of all methods on the Community and Erdos & Renyi datasets. It is obvious that DuSEGO-EGNN can outperform all the baseline methods across these two datasets, supporting the general effectiveness of DuSEGO. It is also worth noting that DuSEGO-EGNN utilizes EGNN as its backbone, the second-order information in DuSEGO-EGNN helps EGNN diminish the % Error and increase the F1 score on these two datasets. Additionally, the experimental results demonstrate that DuSEGO-EGNN remains E(n)-equivariant to the noise distribution, provides the best reconstruction of the adjacency matrix ${A}_{ij}$ on a large-scale graph dataset, and enhances performance when embedding graphs into a continuous space as a set of nodes in dimension n.
Then, we investigate the performance of DuSEGO-EGNN as the number of layers increases (up to 16) compared to EGNN, as shown in Table 4. We observe that increasing depth worsens the F1 score and % Error for EGNN. Additionally, when the number of layers reaches 8 on the Community dataset and 12 on the Erdos & Renyi dataset, gradient explosion occurs, causing EGNN to produce infinite values. These results confirm that gradient explosion happens as EGNN depth increases. However, in contrast to EGNN, DuSEGO-EGNN remains stable on both the Erdos & Renyi and Community datasets, even with up to 16 layers, supporting our hypothesis that gradient explosion is mitigated when EGNN is implemented within DuSEGO.

### v3

@@ 5. DuSEGO Properties Analysis
[…]
Proof:
From (rusch2022graph, ), we have:
@@
Table 1. Mean squared error $\left(\times 10^{-2}\right)$ for the future coordinate estimation under different timesteps in the N-body system experiment, and forward time in seconds for a batch size of 100 samples running on NVIDIA A100 GPUs. Results averaged across five runs. We report both the mean and the standard deviation of MSE in the table. The best performance is bold.
Methods MSE
Forward
Time/ms
@@
6. Experiments
6.1. Modeling a Dynamic System
In this section, we evaluate DuSEGO on one simulated N-body system named charged particles, where we aim to forecast the coordinate of particles across different time steps (satorras2021n, ).
6.1.1. Dataset
Building on the work of (fuchs2020se, ), (satorras2021n, ) extended the Charged Particles N-body experiment introduced in (kipf2018neural, ), where the system consists of five particles, each defined by charge, coordinate, and velocity in three-dimensional space. In this setup, particles in each graph are fully connected, with the initial velocity norm included as an additional node feature. Using the N-body system code from (satorras2021n, ), we generated a new dataset for the charged N-body system across 1,000, 1,500, and 2,000 timesteps. The initial coordinates of particles were sampled from a Gaussian distribution with $\mu=0$ and $\sigma=0.5$, and the initial velocity was represented as a random vector with a norm of 0.5. The dataset is split into training, validation, and testing sets, containing 10,000, 2,000, and 2,000 samples, respectively.
@@ 6.1.1. Dataset
6.1.2. Baselines.
Two types of baseline methods are deployed for fair comparisons: (1) Methods that incorporate first-order information, including the Linear model, Graph Neural Networks (GNN), SE(3) Transformer (fuchs2020se, ), Tensor Field Networks (TFN) (thomas2018tensor, ), Radial Field (RF) (kohler2019equivariant, ), Equivariant Graph Neural Networks (EGNN) (satorras2021n, ), and Steerable E(3) Equivariant Graph Neural Networks (SEGNN) (brandstetter2021geometric, ); (2) Models that incorporate second-order information, including GraphCon (rusch2022graph, ), Graph Mechanics Network (GMN) (huang2021equivariant, ), and Second-order Equivariant Graph Neural Ordinary Differential Equation (SEGNO) (liu2024segno, ).
6.1.3. Experimental Setting.
We build on the experimental framework introduced in (liu2024segno, ), aiming to estimate the coordinates of all particles after 1,000, 1,500, and 2,000 timesteps. We re-evaluate all baseline methods under identical experimental settings on NVIDIA A100 GPUs, gathering results across five runs to ensure fair comparisons. The experimental setup is as follows: Adam optimizer with a learning rate of $5\times 10^{-4}$, batch size of 100, 10,000 epochs, hidden embedding dimension of 64, weight decay of $1\times 10^{-12}$, and four GNN layers. Specifically, for the GMN method, we set the numbers of isolated particles, sticks, and hinges to $(5,0,0)$. For the SE(3)-transformer and TFN, we set the representation degrees to 3 and 2, respectively. An early stopping criterion is applied, halting training if no improvement in validation loss is observed within 50 epochs.
6.1.4. Implementation details.
@@
(b) Dirichlet energy $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node coordinates.
Figure 4. Dirichlet energy $\mathbf{E}\left(\mathbf{H}^{n}\right)$ of node features $\mathbf{H}^{n}$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ of node coordinates $\mathbf{X}^{n}$ propagated through a EGNN and DuSEGO-EGNN for two different values of $\alpha=0,1$ and $\gamma_{1}=\gamma_{2}=1$ in Eq. 8 and Eq. 9.
@@
The results for all baseline methods, DuSEGO-EGNN, and DuSEGO-SEGNN on the N-body system dataset are shown in Table 1. We observe that DuSEGO significantly enhances the performance of EGNN and SEGNN, with DuSEGO-SEGNN outperforming all other baseline methods and DuSEGO-EGNN under the same experimental conditions. Specifically, DuSEGO-EGNN and DuSEGO-SEGNN reduce the mean squared error (MSE) of EGNN and SEGNN by an average of 7.73% and 6.65%, respectively, across the three timesteps, indicating substantial improvement. Notably, the relative improvement for DuSEGO-EGNN increases from 5.86% at 1,000 timesteps to 13.13% at 1,500 timesteps, highlighting the strong generalization ability of DuSEGO. Additionally, Fig. 2 presents a visualization of GMN, EGNN, SEGNN, GMN, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN in estimating particle coordinates for the N-body system experiment at 1,000 timesteps, with MSE values between predictions and ground truths reported. This visualization further shows that DuSEGO-SEGNN and DuSEGO-EGNN outperform EGNN and SEGNN, respectively, aligning with the performance results in Table 1.
@@
It is crucial to verify that the physical principles identified from a limited number of experiments are sufficiently general to explain phenomena on a broader scale. To this end, we assess the performance of EGNN, SEGNN, GMN, GraphCon, SEGNO, DuSEGO-EGNN, and DuSEGO-SEGNN across various training set sizes, ranging from 500 to 3,000 samples in the N-body system dataset. In this evaluation, the timesteps of particle coordinates are set to 1,000. As shown in Fig. 3, DuSEGO-EGNN and DuSEGO-SEGNN, similar to the GMN method, demonstrate more stable performance than other baseline methods, emphasizing their data efficiency. This result suggests that data efficiency, a notable strength of DuSEGO, is essential for effectively learning to model physical systems.
Table 2. Training time (in seconds) per epoch for DuSEGO-EGNN and DuSEGO-SEGNN on the N-body system with 1000 timesteps under different layers.
# Layers DuSEGO-EGNN DuSEGO-SEGNN
4 1.0433 4.0484
8 1.9295 7.7215
12 2.7909 11.7355
16 3.6658 15.5262
@@
To further investigate the over-smoothing issue in EGNN, we implement EGNN and DuSEGO-EGNN with two different values of $\alpha=0,1$ and set $\gamma_{1}=\gamma_{2}=1$ in Eq. 8 and Eq. 9 on the N-body system dataset at 1,000 timesteps. As shown in Fig. 4, the over-smoothing problem in EGNN becomes apparent as the number of EGNN layers increases. The differences between node features, such as $\mathbf{h}_{i}^{n}$ and $\mathbf{h}_{j}^{n}$, as well as node coordinates, like $\mathbf{x}_{i}^{n}$ and $\mathbf{x}_{j}^{n}$, gradually decrease, indicating that node features and coordinates are converging toward a constant value. However, with the integration of DuSEGO into EGNN, the over-smoothing issue is effectively mitigated, as $\mathbf{E}\left(\mathbf{H}^{n}\right)$ and $\mathbf{E}\left(\mathbf{X}^{n}\right)$ remain stable within a reasonable range, even as the number of EGNN layers increases.
@@
Additionally, we compare the performance of EGNN and DuSEGO-EGNN as the number of EGNN layers increases (up to 16). As shown in Table 3, the MSE for EGNN decreases from layer 4 to layer 8. However, at 12 layers, the gradient vanishing issue emerges, preventing EGNN from producing accurate results. This confirms that gradient vanishing becomes a limitation as the depth of EGNN increases. In contrast, the MSE for DuSEGO-EGNN continues to decrease with additional layers across different timesteps, and DuSEGO-EGNN maintains stable performance, supporting our hypothesis that integrating EGNN within DuSEGO helps mitigate the gradient vanishing problem. We also measured the training time per epoch for DuSEGO-EGNN and DuSEGO-SEGNN as the number of GNN layers increased, with timesteps fixed at 1000. As shown in Table 2, training time grows nearly linearly with network layers, consistent with our theoretical analysis in Section 4.3. Notably, DuSEGO-SEGNN shows much higher computational costs per layer than DuSEGO-EGNN, due to the intensive tensor operations in the SEGNN architecture. These results confirm both the scalability and the strong dependence of runtime on the chosen backbone.
@@
Table 3. Mean squared error $\left(\times 10^{-3}\right)$ for the future coordinate estimation under different timesteps in the N-body system experiment between EGNN and DuSEGO-EGNN under different layers. Results averaged across 5 runs.
@@ 6.2. Graph Autoencoder
Subsequently, we compare DuSEGO with the state-of-art baselines for the Graph Autoencoder task, where we aim to estimate the adjacency matrix ${A}_{ij}$ of a graph $\mathcal{G}$ generated from the graph autoencoder and minimize the binary cross entropy between the estimated adjacency matrix $\hat{A}_{ij}$ and the ground truth $A_{ij}$.
@@
Two widely used datasets are deployed in this section: (1) Community  (kipf2016variational, ) dataset, generated using the Erdos-Renyi model (erdds1959random, ), consists of random graphs where nodes are connected with a predefined probability, capturing the diversity in the density and structure of various communities. This model, denoted as $G(n,p)$, generates networks by independently linking pairs of nodes with probability $p$; (2) Erdos & Renyi (janson2011random, ) dataset, based on the foundational random graph generation models (janson2011random, ; you2018graphrnn, ) and also represented as $G(n,p)$, includes a collection of graphs that exhibit the randomness and variability of network structures.
Following (satorras2021n, ), we generate the Community and Erdos & Renyi datasets (kohler2019equivariant, ; kipf2018neural, ) using the code provided in (kipf2018neural, ). Specifically, we sample 50,000 graphs for training, 5,000 for validation, and 5,000 for testing, where we set $p=0.1$ and $6\leq n\leq 10$ for Erdos&Renyi, and $p=0.15$ and $6\leq n\leq 11$ for Community, respectively. Additionally, 20 communities are divided in the Community dataset, and each graph is defined as an adjacency matrix $A\in\{0,1\}^{M\times M}$ in these two datasets.
@@ 6.2.2. Baselines
Two types of baseline methods are deployed for comparison: (1) Methods that contains first-order information including GNN, GNN-Noise, RF (kohler2019equivariant, ), EGNN (satorras2021n, ); (2) Methods that contains second-order information including GraphCon (rusch2022graph, ), GMN (huang2021equivariant, ) and SEGNO (liu2024segno, ).
@@
The noise generated from $\mathcal{N}(\mathbf{0},\sigma\mathbf{I})\in\mathbb{R}^{M\times n}$ will be used as input for all baseline methods. The output will then be employed to embed the graph and fed into the graph decoder to generate an estimated adjacency matrix, $\hat{A}_{ij}$. We rerun all baseline methods under the same experimental settings on NVIDIA A100 GPUs and collect the results across five runs for a fair comparison. The experimental settings are as follows: Adam optimizer with learning rate $1\times 10^{-3}$}, the number of epochs 1,000, batch size 256, hidden embedding dim 64, weight decay $1\times 10^{-12}$}, layer number 4, and dimensions for the embedding space 8. Velocity is not considered in the Graph Autoencoder task, which is set to None in the EGNN, GMN, SEGNO, and DuSEGO-EGNN.
@@
We build upon the experimental setting introduced in (satorras2021n, ) and employ EGNN as the backbone of DuSEGO, named DuSEGO-EGNN. For this experiment, we report the F1 score and % Error for all used models, where % Error is defined as the percentage of incorrectly predicted edges relative to the total number of potential edges. EGNN is implemented based on the implementation details provided in (satorras2021n, ), which is used as the coupling function for message passing (rusch2022graph, ). Then, we use the ODE solvers provided in the torchdiffeq library for solving Eq. 8 and Eq. 9.
@@
Table 4 depicts the overall results of all methods on the Community and Erdos & Renyi datasets. It is evident that our DuSEGO-EGNN can outperform all other methods across these two datasets, supporting the general effectiveness of DuSEGO. It is also worth noting that DuSEGO-EGNN utilizes EGNN as its backbone; the second-order information in DuSEGO-EGNN helps EGNN diminish the % Error and increase the F1 score on these two datasets. Additionally, the experimental results demonstrate that DuSEGO-EGNN remains E(n)-equivariant to the noise distribution, provides the best reconstruction of the adjacency matrix ${A}_{ij}$ on a large-scale graph dataset, and enhances performance when embedding graphs into a continuous space as a set of nodes in dimension n.
Then, we investigate the performance of DuSEGO-EGNN when the number of layers increases (up to 16) compared to EGNN, as shown in Table 5. We observe that a higher layer can lead to a lower F1 score and higher % Error for the baseline EGNN model. Additionally, when the number of layers reaches eight on the Community dataset and 12 on the Erdos & Renyi dataset, gradient explosion occurs, causing EGNN to produce infinite values. This phenomenon demonstrates that the gradient explosion occurs as the number of EGNN layers increases. On the contrary, our proposed DuSEGO-EGNN remains stable on both the Erdos & Renyi and Community datasets, even with up to 16 layers, supporting our hypothesis that gradient explosion is mitigated when EGNN is implemented within DuSEGO.